OpenStudy (anonymous):

Help Please? Factor Completely: 2x2 − 32 Prime 2(x2 − 16) 2(x + 4)(x + 4) 2(x + 4)(x − 4)

OpenStudy (aroub):

2(x + 4)(x − 4)

OpenStudy (anonymous):

thanks... how do you figure that out?

OpenStudy (aroub):

ok first b/c u have 2 in common so put the two out

OpenStudy (aroub):

then u get 2(x^2-16)

OpenStudy (aroub):

then 2(x-4)(x+4)

OpenStudy (aroub):

lol sry ik im not gd at explaining..

OpenStudy (anonymous):

ok... i think i get it now... thanks!!! :)

OpenStudy (anonymous):

2 is common so take it out \[2(x^2-16)\] \[x^2-y^2=(x+y)(x-y)\] so,\[2[(x)^2-(4)^2]=2(x+4)(x-4)\]